Calculate ΔG0 for the following reactions at 25oC.

(a) N2(g) + O2(g) → 2NO(g)

ΔG0 = kJ/mol

(b) H2O(l) → H2O(g)

ΔG0 =    kJ/mol

(c) 2C2H2(g) + 5O2(g) → 4CO2(g) + 2H2O(l)

ΔG0 = kJ/mol
Answer #1

ΔG0 rxn=\sumn \DeltaGfo (products)-\summ\DeltaGfo (products)

\DeltaGfo-std gibbs energy of formation

n,m=coefficients of reactants

\DeltaGfo (N2)=0

\DeltaGfo (O2)=0

\DeltaGfo (NO)=87.6 KJ/mol

ΔG0 rxn=2*87.6 KJ/mol-0=175.2kj/mol

b) \DeltaGfo (H2O L)=-237.14 KJ/MOL

\DeltaGfo (H2O GAS)=0

ΔG0 rxn=0--237.14 KJ/MOL=-237.14 KJ/MOL

C)\DeltaGfo (C2H2))=209 KJ/mol

\DeltaGfo (o2)=0

\DeltaGfo (CO2)=-394.39kj/mol

\DeltaGfo (H2O L)=-237.14 KJ/MOL

\DeltaG 0 rxn=[4*(-394.39)+2*(-237.14)]-[2*209+0]=-2051.84-418=-2469.84 kj/mol